### v2

@@ 2.3 Model evaluation
The Area Under the Precision-Recall Curve (PR AUC) measures the trade-off between precision and recall for different threshold values. Precision is defined as the proportion of true positives among all positive predictions, while recall (or sensitivity) is the proportion of true positives among all actual positives. PR AUC is particularly useful for imbalanced datasets, where the number of negatives far exceeds the number of positives. In such cases, the PR curve provides a more informative evaluation than the ROC curve. The PR AUC is computed using the ‘average_precision_score‘ function in scikit-learn, which calculates the weighted mean of precision scores at each threshold:
@@ 3.1 Comparison Experiments
After normalizing the datasets, we trained each task using esm_ratio values of 0.9, 0.95, and 1, along with lambda_ values ranging from 0.9 to 1 (with a step size of 0.1). The model with the best performance, determined based on ACC, AUC, and PR metrics, was selected as the optimal configuration for this experiment.
We accessed the UniDL4BioPep model’s server to obtain results on the same benchmark dataset. The experiments showed that, compared with the UniDL4BioPep model, our model’s ACC, AUC, and PR were on average higher by 0.3%, 0.7%, and 0.8%. Furthermore, the FN and FP were reduced by 76 and 14 samples. On the ten datasets where the model performed best, our model’s ACC increased by 2%, AUC increased by 1.67%, and PR increased by 1.87%, while the FN and FP were reduced by 71 and 67 samples, respectively.The comparison of various metrics on different datasets among PDeepPP, UniDL4BioPep, and the tools compared in UniDL4BioPep is shown in SuppTable 2.
Specifically, compared with the UniDL4BioPep model, our model shows significant advantages in identification accuracy (ACC), indicating more stable and efficient performance in the overall classification task. Additionally, the improvement in AUC suggests that our model has enhanced ability to discriminate between positive and negative samples, better capturing complex patterns in the data. The improvement in the PR curve demonstrates that our model is more robust in handling imbalanced data, particularly achieving better predictive performance for minority classes (such as positive samples).The ROC and PR curves and confusion matrices for all tasks are shown in Figure 2 and SuppFig 1.
In terms of false negatives (FN) and false positives (FP), our model also demonstrates superior performance. The reduction in FN and FP indicates that our model makes more cautious identification decisions, particularly in the test set’s positive and negative samples, effectively reducing the number of misclassifications. Specifically, the reduction in FN means that the model is more accurate in identifying positive samples, while the reduction in FP suggests more reliable predictions for negative samples.
@@
In the comparison with the MusiteDeep model, we also used the same benchmark dataset. The results indicate that, compared with the MusiteDeep model, our model’s ACC, AUC, and PR were on average higher by 3.3%, 0.1%, and 1.9%. Additionally, FN increased by 1,111, while FP decreased by 4,754 samples. On the six datasets where the model performed best, our model’s ACC increased by 0.8%, AUC increased by 5.0%, and PR increased by 1.6%.Even if FN increased by 3,096, FP decreased by 21,436. This indicates that the model demonstrates a significantly stronger capability in identifying negative samples on the highly imbalanced dataset, but it also leads to a deficiency in recognizing positive samples.The comparison of various metrics on different datasets between PDeepPP and MusiteDeep is shown in SuppTable 3.
@@
Figure 3: The top six tasks with the best performance in comparison with the MusiteDeep model. The upper section shows the AUC curves and confusion matrix combinations for each task. In the confusion matrices, the numbers in parentheses next to the values represent the difference in the count of this metric compared to PDeepPP for the MusiteDeep model. The color intensity of each block in the matrix indicates the proportion of that class relative to the total negative/positive samples. The lower section displays the UMAP plots for each dataset after feature extraction by PDeepPP, with the site distribution of these six datasets shown in the bottom right. The label order is: (A) Phosphorylation_Y, (B) N-linked-glycosylation_N, (C) N6-acetyllysine_K, (D) Methylation_K, (E) Ubiquitin_K, (F) Hydroxyproline_K.
In terms of classification accuracy (ACC), our model outperforms the MusiteDeep model, indicating higher stability and effectiveness in the overall classification task. The improvements in AUC and PR further suggest that our model has enhanced ability to distinguish between positive and negative samples, better adapting to different data distributions. Moreover, the improvement in the PR curve also demonstrates the advantage of our model in handling imbalanced data, particularly showing stronger sensitivity and accuracy in dealing with positive samples.The ROC and PR curves and confusion matrices for all tasks are shown in Figure 3 and SuppFig 2.
Similarly, in terms of the reduction in FN and FP, our model outperforms the MusiteDeep model. The reduction in FN and FP means that our model is more accurate in identifying positive samples and predicting negative samples, effectively avoiding misclassifications. Particularly in predicting negative samples, our model demonstrates significant advantages in terms of FP, which indicates that the risk of misclassifying negative samples as positive in practical applications will be greatly reduced.
@@ 3.2 Feature Extraction Analysis
3.2.1 1. Introduction to ESM Feature Extraction
ESM (Evolutionary Scale Modeling) is a large-scale protein language model based on the Transformer architecture and trained with masked language modeling on massive protein sequences from the UniRef database. By leveraging relative positional embeddings and multi-head self-attention, ESM captures evolutionary information and models amino acid interactions to generate high-dimensional representations that encode secondary and tertiary structural features. With up to 15 billion parameters, ESM-2 achieves significant improvements in perplexity and structure prediction accuracy, demonstrating strong correlations between language model understanding and atomic-level structure prediction. Its embeddings serve as inputs for ESMFold, enabling rapid and accurate end-to-end protein structure prediction without relying on MSAs, offering a computationally efficient alternative to traditional methods.
In this study, due to ESM’s specificity for proteins, ESM is utilized as the primary feature extraction tool, and the generated protein sequence embeddings are further applied to downstream analysis tasks. The core idea of feature extraction relies on the deep understanding of sequences by the ESM model, thereby providing biologically meaningful representations for downstream tasks. We adopt the residual connection strategy by adding an embedding layer based on linear transformations and perform weighted feature extraction to enhance representation capacity, optimizing the performance of downstream tasks.
3.2.2 2. UMAP Visualization of Model Features
@@
In this study, UMAP (Uniform Manifold Approximation and Projection) is utilized to visualize the high-dimensional embeddings extracted before the prediction moduleUMAP ; umap2 . (UMAP is a nonlinear dimensionality reduction technique that projects high-dimensional data into a two-dimensional space, enabling intuitive observation of data distribution and clustering patterns.) These embeddings are derived from the combined outputs of the TransLinear layer and the PosCNN layer, averaged along the sequence dimension. From the UMAP visualization, it can be observed that the features generated by the model exhibit effective class separation after dimensionality reduction, with distinct clusters formed in the reduced feature space for different categories.
By visualizing with UMAP, researchers can intuitively observe the clustering patterns and inter-class differences of different protein features, providing important insights into the interpretability of the model and the relationships between features.
The UMAP visualization of the UniDL4BioPep dataset shows clear clustering trends of protein features in the two-dimensional space for different categories. This indicates that the features extracted by the model have strong discriminative power. The specific analysis is as follows:
3.2.3 3. Analysis of UniDL4BioPep UMAP Visualization
(Figure 2 and SuppFig 1)
@@
3.2.4 4. Analysis of MusiteDeep UMAP Visualization
(Figure 3 and SuppFig 2)
@@
Figure 4: The best-performing datasets in comparison with the UniDL4BioPep and MusiteDeep models (10 datasets for UniDL4BioPep and 6 for MusiteDeep) and the average values of metrics across all datasets.
3.3 Ablation Study
In this subsection, we conduct an ablation study to evaluate the contribution of key modules in our modelablation . By systematically removing specific components, we analyze the impact on performance using ROC/PR curves, confusion matrices, and UMAP visualizations.
Dataset Selection
For dataset selection, we first chose four datasets from the UniDL4BioPep and MusiteDeep collections based on dataset size, the ratio of positive to negative sites, and prediction performance. These included Anticancer_main and Antiviral from UniDL4BioPep, and N6-acetyllysine_K and Ubiquitin_K from MusiteDeep, which served as the first batch of datasets for the ablation study. To explore the model’s performance on tasks with additional classification targets, similar post-translational modification (PTM) types with emphasis on different amino acids, and tasks where both PTM type and target amino acid are the same, we selected several recent state-of-the-art (SOTA) datasets from recent literature. These datasets include methylation-Gmethylation-G , methylation-R (distinguished from the corresponding task in MusiteDeep by the lowercase letter)methylation-R , Ubiquitin_K* (the asterisk is used to differentiate it from the corresponding task in MusiteDeep)ubiquitin , and Crotonylation_Kkcr . The goal of this selection is to verify the model’s robustness and completeness by testing it on diverse datasets while ensuring that each module contributes positively to the overall model performance.The specific sample sizes and the distribution of negative and positive samples across datasets (especially for imbalanced datasets) are shown in SuppTable 1 and Figure 5, respectively.
Experimental Setup
We design several ablation experiments:
  • Full model (baseline).
  • Removing pre-trained embeddings (sequence or structure).
  • Replacing the TIM loss function with a standard cross-entropy loss.
  • Removing the PosCNN(&PosEncoding) or
    TransLinear(&First attention layer) module.

### v3

@@ 2.3 Model evaluation
The Area Under the Precision-Recall Curve (PR AUC) measures the trade-off between precision and recall for different threshold values. Precision is defined as the proportion of true positives among all positive identifications, while recall (or sensitivity) is the proportion of true positives among all actual positives. PR AUC is particularly useful for imbalanced datasets, where the number of negatives far exceeds the number of positives. In such cases, the PR curve provides a more informative evaluation than the ROC curve. The PR AUC is computed using the ‘average_precision_score‘ function in scikit-learn, which calculates the weighted mean of precision scores at each threshold:
@@ 3.1 Comparison Experiments
We conducted hyperparameter optimization over a range of esm ratio and lambda values to identify optimal configurations for each task. To ensure consistent training across all datasets, we applied standardized data preprocessing and normalization procedures. The detailed tuning strategies are provided in the Supplementary Material (see Section: Hyperparameter Optimization Details). Results from the UniDL4BioPep model server on the same benchmark datasets show that PDeepPP achieves 0.3% higher average accuracy (ACC), 0.7% higher ROC AUC, and 0.8% higher PR AUC compared to UniDL4BioPep. Specifically, false negatives (FN) and false positives (FP) are reduced by 76 and 14 samples, respectively this highlights the model’s ability to minimize both false predictions through its hybrid feature fusion strategy. On the top 10 performing datasets, PDeepPP demonstrates 2% higher ACC, 1.67% higher AUC, and 1.87% higher PR with 71 fewer FN and 67 fewer FP (SuppTable 2), evidencing its robustness across diverse biological tasks.
@@
Figure 3: The top six tasks with the best performance in comparison with the MusiteDeep and PhosF3C models. The upper section shows the AUC curves and confusion matrix combinations for each task. In the confusion matrices, the numbers in parentheses next to the values represent the difference in the count of this metric compared to PDeepPP for both the MusiteDeep and PhosF3C models (shown as MusiteDeep,PhosF3C). The color intensity of each block in the matrix indicates the proportion of that class relative to the total negative/positive samples. The lower section displays the UMAP plots for each dataset after feature extraction by PDeepPP, with the site distribution of these six datasets shown in the bottom right. The label order is: (A) Phosphorylation_Y, (B) N-linked-glycosylation_N, (C) N6-acetyllysine_K, (D) Methylation_K, (E) Ubiquitin_K, (F) Hydroxyproline_K.
PDeepPP consistently outperforms MusiteDeep and PhosF3C across PTM tasks, with the architecture’s parallel CNN-Transformer design enabling it to capture both local modification motifs and global sequence context. For instance, in Hydroxyproline K identification, PDeepPP achieves an ACC of 0.9885, markedly higher than MusiteDeep’s 0.9412 and PhosF3C’s 0.9476 (SuppTable 3). When compared to MusiteDeep, PDeepPP shows 4.51% higher average ACC and 6.2% higher PR AUC. Across all PTM datasets, PDeepPP maintains an average ACC of 0.9847, surpassing MusiteDeep (0.9396), a testament to the effectiveness of ESM-2 pretrained embeddings in encoding structural knowledge. In the comparison with the MusiteDeep model on average, our model’s ACCand PR were higher by 3.3% and 1.9% respectively. Additionally, FN increased by 1,111, while FP decreased by 4,754 samples. On the six datasets where the model performed best, our model’s ACC increased by 0.8%, AUC increased by 5.0%, and PR increased by 1.6%.Even if FN increased by 3,096, FP decreased by 21,436. The comparison of various metrics on different datasets between PDeepPP and MusiteDeep is shown in SuppTable 3.
Similarly, the comparison against PhosF3C highlights the strengths of PDeepPP. Across all PTM datasets, PDeepPP achieves a 3.33% higher average ACC, a 3.48% higher ROC AUC, and a significant 11.92% higher PR AUC. In the specific task of Phosphorylation Y identification, PDeepPP not only demonstrates a clear advantage in overall accuracy (ACC: 0.9944 vs. 0.9486) but also shows a substantially higher Matthews Correlation Coefficient (MCC: 0.9654 vs. 0.7194), indicating a much better balance between sensitivity and specificity. This superior performance is consistent across the board, with PDeepPP showing an average ACC of 0.9847 and MCC of 0.9012, compared to PhosF3C’s average ACC of 0.9514 and MCC of 0.7542 (SuppTable 3). These results collectively validate PDeepPP’s ability to balance computational efficiency with state-of-the-art predictive accuracy.
The ROC and PR curves and confusion matrices for all tasks are shown in Figure 3 and SuppFig 2. For a comparison of the top 10 datasets from UniDL, the top 6 datasets from MusiteDeep, and the average metrics across all datasets, please refer to Figure 4.
The ROC and PR curves and confusion matrices for all tasks are shown in Figure 3 and SuppFig 2.
@@ 3.2 Feature Extraction Analysis
In this study, UMAP (Uniform Manifold Approximation and Projection) is utilized to visualize the high-dimensional embeddings extracted before the identification moduleUMAP ; umap2 . (UMAP is a nonlinear dimensionality reduction technique that projects high-dimensional data into a two-dimensional space, enabling intuitive observation of data distribution and clustering patterns.) These embeddings are derived from the combined outputs of the TransLinear module and the PosCNN module, averaged along the sequence dimension. From the UMAP visualization, distinct clusters are formed in the reduced feature space for different categories. UMAP visualizations of PDeepPP demonstrate distinct cluster separations and sharp boundaries for functional peptides and PTM sites, outperforming UniDL4BioPep, MusiteDeep, and PhosF3C. The visualizations highlight well-defined clusters for various peptide classes and modification types, reflecting the model’s effectiveness in capturing discriminative sequence features and contextual relationships. A comprehensive analysis of the UMAP results—including cluster distribution patterns, boundary sharpness, and outlier behavior in comparative experiments with UniDL4BioPep and MusiteDeep—is provided in the Supplementary Material (see Section: Detailed UMAP Visualization Analysis for Comparison Experiments).
Figure 4: Performance comparison of the best-performing datasets previously shown in Figure 2 (10 datasets for UniDL4BioPep) and Figure 3 (6 datasets for MusiteDeep and PhosF3C), along with the average values of metrics across all datasets. The left panel shows the comparison between PDeepPP and UniDL4BioPep, while the right panel displays the comparison among PDeepPP, MusiteDeep, and PhosF3C. These datasets represent the top-performing cases detailed in the UMAP visualizations and confusion matrices of the preceding figures. The solid lines represent the performance on best datasets, while the dashed lines indicate the average performance across all datasets.
3.3 Ablation Study
To rigorously validate the superiority of our hybrid design philosophy, we conduct comprehensive ablation experiments that systematically evaluate each component’s contribution to the overall framework performance. Our ablation strategy is specifically designed to demonstrate why the synergistic combination of pretrained embeddings, parallel neural architectures, and specialized loss functions outperforms individual components or simpler design alternatives.ablation . By systematically removing specific components, we analyze the impact on performance using ROC/PR curves, confusion matrices, and UMAP visualizations.
Dataset Selection
For dataset selection, we first chose four datasets from the UniDL4BioPep and MusiteDeep collections based on dataset size, the ratio of positive to negative sites, and identification performance. These included Anticancer_main and Antiviral from UniDL4BioPep, and N6-acetyllysine_K and Ubiquitin_K from MusiteDeep, which served as the first batch of datasets for the ablation study. To explore the model’s performance on tasks with additional classification targets, similar post-translational modification (PTM) types with emphasis on different amino acids, and tasks where both PTM type and target amino acid are the same, we selected several recent state-of-the-art (SOTA) datasets from recent literature. These datasets include methylation-Gmethylation-G , methylation-R (distinguished from the corresponding task in MusiteDeep by the lowercase letter)methylation-R , Ubiquitin_K* (the asterisk is utilized to differentiate it from the corresponding task in MusiteDeep)ubiquitin , and Crotonylation_Kkcr . The specific sample sizes and the distribution of negative and positive samples across datasets (especially for imbalanced datasets) are shown in SuppTable 1 and Figure 5, respectively.
Experimental Setup
We conducted targeted ablation experiments to validate the core principles underlying our hybrid design. Each experiment is designed to test a specific hypothesis: (1) pretrained knowledge enhances task-specific learning; (2) parallel processing of global and local features outperforms single-path architectures; and (3) information-maximization objectives improve performance on imbalanced biological datasets. The complete experimental setup is described in the Supplementary Material (see Section: Detailed Ablation Experimental Setup).
Hybrid Design Validation Framework
Our ablation methodology specifically tests three core hypotheses underlying the hybrid design:
Hypothesis 1 - Complementary Feature Representation: The combination of ESM-2 (capturing evolutionary patterns) and BaseEmbedding (learning task-specific motifs) provides complementary information that neither component achieves alone.
Hypothesis 2 - Parallel Processing Superiority: Simultaneous extraction of global dependencies (via Transformer) and local motifs (via CNN) followed by feature fusion outperforms sequential or single-pathway processing.
Hypothesis 3 - Information-Theoretic Optimization: TIM loss maximizes mutual information between input sequences and labels, providing superior handling of class imbalance compared to standard cross-entropy approaches.
Results and Analysis: Hybrid Design Superiority
Evaluation metrics
The hybrid architecture model (PDeepPP) demonstrates strong performance across all evaluation metrics, achieving a median Area Under the Curve (AUC) of 0.991, an accuracy (ACC) of 0.8349, and a balanced accuracy (BACC) of 0.8282, highlighting the effectiveness of its design. The model’s primary advantage lies in the synergistic effect produced by its parallel CNN and Transformer modules. Ablation studies show these modules are functionally complementary: removing the CNN module impairs local feature extraction, causing sensitivity (SN) to drop to 0.7915 from a higher baseline and increasing false negatives (FN). Conversely, removing the Transformer (and its attention mechanism) weakens global context modeling, leading to more false positives (FP). This complementary error pattern confirms that the hybrid design effectively balances local and global information, which is critical for precise peptide identification.
Analysis of component contributions reveals that while all modules positively impact performance, some are of key importance. The modules handling positional information—Positional Encoding (PosEncoding) and the Positional CNN (PosCNN)—are the most significant contributors. Removing them causes the Matthews Correlation Coefficient (MCC) to decrease substantially from a baseline of 0.8265 to 0.6511 and 0.5503, respectively, resulting in a significant performance loss. Similarly, the sequence embedding is a key component of the model; its removal leads to a substantial 9.7% drop in AUC. In contrast, other embedding modules have a lesser impact. In summary, the model’s strong overall performance originates from a complete and synergistic system where every component plays an important role, with the precise encoding of sequence and positional information being central to its success.
UMAP Analysis of Ablation Study A comprehensive analysis of the UMAP visualization results across the eight datasets suggests the following:
  1. 1.
    Contribution of the Hybrid Architecture: Removing either the TransLinear (global features) or PosCNN (local features) module generally has a significant negative impact on the model’s feature representation capability. This suggests that the model’s performance benefits from the combination of both global and local feature extractors.
  2. 2.
    Benefit of the Information-Theoretic Loss: Replacing the TIM loss with a standard cross-entropy loss (w/o loss) tends to reduce model performance, particularly on class-imbalanced PTM datasets. This is visualized as less defined cluster boundaries and increased class overlap, highlighting the benefit of the TIM loss in learning discriminative features.
  3. 3.
    Incremental Contributions: Removing the task-adaptive BaseEmbedding or other sub-components (e.g., attention, positional encoding) also affects model performance, though often to a lesser extent than removing a full architectural branch or the TIM loss. This indicates that each part of the model provides a valuable contribution to the final feature representation. It is worth noting, however, that even the full PDeepPP model does not achieve perfect separation on all datasets, suggesting that for the most challenging cases, sequence information alone may have its limitations.
A detailed analysis of UMAP visualizations and corresponding performance metrics from the ablation study is presented in the Supplementary Material (see Section: UMAP Analysis of Ablation Study).
4 Conclusion
The PDeepPP framework represents a significant advancement in peptide and PTM identification, demonstrating the capability of a unified deep learning approach to address complex bioinformatics challenges. Its success stems from an innovative integration of a pretrained protein language model (ESM-2) with a parallel neural network architecture that combines Transformers and CNNs. This hybrid design proves highly effective, allowing the model to simultaneously capture both global, long-range dependencies and local, conserved sequence motifs, thereby creating a comprehensive and hierarchical feature representation for diverse biological tasks.
Our experimental results show that PDeepPP consistently outperforms state-of-the-art models across a wide array of tasks. In a comprehensive benchmark of 33 datasets, it achieves superior performance in 25 tasks, attaining 0.9726 accuracy in antimicrobial peptide identification and 0.9984 accuracy in phosphorylation site prediction, while substantially reducing false positives and false negatives. This robust performance is not only a product of its hybrid architecture but also its use of the Transductive Information Maximization (TIM) loss function, which effectively addresses the pervasive issue of class imbalance in biological datasets and enhances the model’s utility in real-world scientific scenarios.
The framework demonstrates remarkable generalization, suggesting its potential as a general-purpose computational platform that can reduce the need for developing bespoke models for each specific identification task. However, we acknowledge its limitations. As a sequence-based model, PDeepPP does not explicitly incorporate three-dimensional structural information, which could further enhance predictive accuracy for PTM sites or functional domains that are heavily dependent on spatial conformation. Furthermore, while visualization techniques such as UMAP offer insights, the inherent ”black-box” nature of deep learning models makes it challenging to fully elucidate the biological mechanisms behind their predictions.
Looking ahead, PDeepPP opens several promising avenues for future research and application. There is great potential in integrating multi-modal data, such as predicted protein structures from models such as AlphaFold, cellular localization data, or gene co-expression networks, to build a more holistic predictive framework. The model could also be extended to support multi-label or multi-task learning, enabling the simultaneous prediction of different PTM types on the same peptide or combining function prediction with other relevant tasks, such as toxicity or stability. A key future goal will be to develop more advanced explainable AI techniques to extract specific, biologically verifiable sequence patterns from the model, moving beyond feature visualization to mechanistic insight. In summary, PDeepPP offers a cutting-edge solution for protein sequence analysis that has the potential to accelerate peptide discovery, PTM functional analysis, and therapeutic strategy development, significantly reducing reliance on costly and time-consuming experimental methods.
5 Data and Code Availability
6 Key Points
  • PDeepPP integrates pretrained language models with a novel hybrid architecture for peptide identification.
  • Achieves state-of-the-art performance across 33 tasks, outperforming existing approaches.
  • Effectively captures both local and global sequence features, validated by advanced analysis.
  • TIM loss and ablation studies demonstrate the model’s effectiveness; code is publicly available.
Supplementary Figure 1.
Comparison of PDeepPP and UniDL4BioPep on ten additional bioactive peptide tasks, with AUC/confusion matrices and UMAP plots.
Supplementary Figure 2.
Comparison of PDeepPP and MusiteDeep on six additional PTM tasks, including AUC/confusion matrices and UMAP plots.
Supplementary Figure 3.
UMAP visualizations for ablated models and PDeepPP across four datasets.
8 Supplementary
8.1 Detailed Architecture Specifications
8.1.1 BaseEmbedding Module Architecture
The BaseEmbedding module utilizes a task-adaptive architecture. It employs a dynamic, learnable layer to map each amino acid into a 128-dimensional space optimized for peptide identification. This representation is then transformed by a specialized linear projection layer, scaling it to 1280 dimensions to align with ESM-2 embeddings. This layer also incorporates learned bias terms to capture task-specific amino acid relationships not present in general protein models. The design enables the model to learn complementary representations that focus on local sequence motifs and functional patterns specific to tasks like PTM site identification, allowing it to capture fine-grained features while maintaining compatibility with the rich contextual information from ESM-2.
8.1.2 TransConv1d Network Architecture
The TransConv1d module combines self-attention mechanisms, a Transformer encoder, and fully connected networks to extract global and local features from input data, making it suitable for processing sequential data or high-dimensional features. Its architecture consists of multiple modules, each with specific hyperparameters to ensure the model’s expressive capability and training stability.
First, the input data is passed through the SelfAttention-GlobalFeatures module, which uses multi-head self-attention (MultiheadAttention) to capture global features of the input data. Specifically, the [embed_dim] is set to the input feature dimension, and [num_heads] is set to 8 to parallelize the processing of different feature patterns. The output of the self-attention is added to the original input via a residual connection and then stabilized with layer normalization (LayerNorm). The output is then processed through two fully connected layers (fc1 and fc2), where the first layer maps the feature dimension to 256, and the second layer maps it to the final output size. Dropout is applied after the first fully connected layer to prevent overfitting.
By combining self-attention mechanisms, the Transformer encoder, and fully connected layers, the TransConv1d module effectively captures the complex dependencies in the input data. The residual connections and layer normalization help maintain data stability and preserve information, thus improving the model’s expressive power and training stability.
8.1.3 PosCNN Network Architecture
The PosCNN module integrates convolutional neural network (CNN) operations with optional positional encoding, designed to extract local features from input data while preserving sequence information, making it suitable for processing sequential data or high-dimensional features. The architecture focuses on capturing local patterns through convolution operations, followed by a fully connected layer for feature mapping.
The overall structure of the PosCNN module is designed to extract local features from the input sequence, enhance these features with positional information (if enabled), and finally map them to the output space using a fully connected layer.
8.2 Hyperparameter Optimization Details
After normalizing the datasets, we trained each task using esm_ratio values of 0.9, 0.95, and 1, along with lambda values ranging from 0.9 to 1 (with a step size of 0.1). The model with the best performance, determined based on ACC, AUC, and PR metrics, was selected as the optimal configuration for this experiment.
8.3 Detailed UMAP Visualization Analysis for Comparison Experiments
8.3.1 Analysis of UniDL4BioPep UMAP Visualization
@@
8.3.2 Analysis of MusiteDeep UMAP Visualization
@@
8.4 Detailed Ablation Experimental Setup
We design targeted ablation experiments to validate our hybrid design principles:
  • Embedding Integration Validation: We remove pretrained embeddings (ESM-2) to demonstrate the necessity of leveraging large-scale protein knowledge, and remove BaseEmbedding to show the importance of task-specific feature learning.
  • Parallel Architecture Validation: We systematically remove the TransLinear module (global feature extraction) and PosCNN module (local feature extraction) to validate that the parallel combination of transformer and CNN architectures is superior to using either component alone.
  • Loss Function Optimization Validation: We replace the TIM loss with standard cross-entropy to demonstrate the effectiveness of our information-theoretic approach for handling imbalanced datasets.
  • Hybrid vs. Sequential Design: We compare our parallel fusion approach against sequential processing alternatives to validate the architectural choice.